Solve each equation for x.
(a) 30 = 80(1 − e^−2x)
(b) 5 × 2^x = 3^ (x−1)
(c) log2(log3 x) = 5
(d) e^{− 1/2 log(x^2−1) }=1/sqrt 3

Answer by Alan3354(69443)

(a) 30 = 80(1 - e^-2x)
1 - e^-2x = 0.375
e^-2x = 0.625
-2x = ln(0.625)
x = -ln(0.625)/2
x =~ 0.2350018
======================
(b) 5 × 2^x = 3^ (x−1)
Not clear. Use * (Shift 8) for multiply, not x.
================
(c) log2(log3 x) = 5
Ambiguous. Add parentheses.
===================
(d) e^{-1/2 log(x^2 - 1) }=1/sqrt 3
Is it log base 10? or ln ?